### v3

@@ Abstract
Due to the dynamic nature of the Semantic Web, version control is necessary to capture time-varying information, particularly for widely used ontologies. Despite the long-standing recognition of ontology versioning (OV) as a crucial component for efficient ontology management, the growing size of ontologies and accumulating errors caused by manual labour overwhelm current OV approaches. In this paper, we propose yet another approach to performing OV using existing ontology matching (OM) techniques and systems. We introduce a unified OM4OV pipeline. From an OM perspective, we reconstruct a new task formulation and measurement for OV tasks. Building upon the prior alignment(s) from OM, we propose a pipeline optimisation method called the cross-reference (CR) mechanism to enhance overall OV performance. We experimentally validate the OM4OV pipeline and the cross-reference mechanism in the OV tested originating from the Ontology Alignment Evaluation Initiative (OAEI) datasets. We also discuss insights into OM used for OV tasks, where some false mappings detected by OV systems are not actually untrue.
@@ 1 Introduction
Ontologies serve as the backbone of the semantic web, providing formal descriptions of shared concepts across various applications [9]. An ontology is not static, and the need for version control arises with its birth. While web data is dynamic, any ontology used must undergo periodic revisions to keep pace with the growth in domain knowledge, modifications to application adaptation, or corrections to the shared conceptualisation [14]. For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” [10]. This may cause undesirable deficiencies in artefacts that conform to or reuse the ontology that is being changed, leading to severe non-compliance and incompatibility issues in downstream tasks.
@@
Ontology versioning (OV) aims to distinguish and recognise changes between different ontology versions. By doing so, data that conforms to the changed ontology, other ontologies that reuse the changed ontology, or software that uses the changed ontology, can apply the correct changes correspondingly [14]. While various methods for OV have been developed, one approach is to extend the ontology itself with internal version information. An ontology can be issued with a unique identifier (e.g. IRI) or a specific version number (e.g. owl:versionInfo) to be distinguished from other versions; each ontology entity may have a new triple to record its current status (e.g. owl:DeprecatedClass and owl:DeprecatedProperty); or every triple could be extended with a 4th dimension populated with a triple timestamp, similar to the RDF-star schema. Alternatively, version information can be recorded in change logs. Change logs can take the form of notebooks, an extensible markup language (XML), or a knowledge graph (KG). However, maintaining version information in the ontology can be time-consuming and labour-intensive. Either extending the current schema or using change logs requires consistent updating over time. In most cases, this process is hand-crafted by the ontology engineer or requires human intervention (e.g. pre-defining a schema or creating a template for the change log). A manual process is more likely to make mistakes and fail to propagate changes to dependent artefacts. Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or has a separate change log. In such cases, current approaches have limited capabilities in detecting incorrect version information or identifying missing version information.
@@
In this study, we investigate a lightweight and fully automatic version control approach for ontologies. We observe that the nature of OV is very similar to that of ontology matching (OM). Both are introduced to address interoperability between ontologies, with ontology entities serving as inputs. While OM is a well-studied problem [19], a unified pipeline that can extend and reuse existing OM techniques and systems for OV tasks would be highly desirable. However, OM techniques and systems cannot be directly applied to OV tasks because OM and OV have some significant differences. (1) The OM input is two distinct ontologies, while the OV input is two different versions of a single ontology. The output of OM is a set of entity mappings, whereas the output of OV is two sets: changed entities and unchanged entities. (2) OM concentrates on similarities between two entities, while OV addresses differences between versions of one entity. OM determines matched entities between two different ontologies, while OV distinguishes add, delete, remain, and update entities between different versions in one ontology. To address these challenges, our goal is to investigate the complementary relationship between these two tasks and to develop a unified approach to shifting OV tasks towards OM tasks. Specifically, our key contributions include:
@@
  • We find that some false mappings detected by OV systems are not actually untrue, but can be caused by a flawed choice of ontology design, an ambiguous “equivalent” relationship, or an inappropriate setting of the similarity threshold.
The remainder of the paper is organised as follows. Section 2 reviews the literature on OV. Section 3 introduces our novel OM4OV pipeline, while Section 4 evaluates the proposed pipeline. We propose our pipeline optimisation method (i.e. cross-reference mechanism) and experimentally evaluate its performance in Section 5. Section 6 and 7 discuss our implications and current limitations. Section 8 concludes the paper.
@@ 2 Related Work
One option is to include version information within the ontology. The Simple HTML Ontology Extensions (SHOE) [12] uses the tag BACKWARD-COMPATIBLE-WITH to record version information. The authors in [13] argue that a carefully-managed version numbering system embedded in the URI of the ontology (and therefore the fully expanded name of entities defined in the ontology) can minimise the impact of adopting updated versions because unchanged entities will be unaffected in practice. These approaches have largely been adopted by the later OWL Web Ontology Language [4], where a set of annotation properties related to version information is defined. These include owl:versionInfo and owl:priorVersion to describe the version number of the ontology, owl:backwardCompatibleWith and owl:incompatibleWith to specify the entity’s compatible or incompatible corresponding entity in the previous version, and owl:DeprecatedClass and owl:DeprecatedProperty to declare the deprecated entities. Later, the ontology language $\tau$OWL [24] was introduced to extend the OWL triple schema to quadruples to represent the versioning of concepts within an ontology. This idea is now incorporated in the new proposals for RDF 1.2, which allow time-varying information to be deduced from a temporal dimension within the quadruple [22].
@@
While current approaches simply record human-generated version information, less attention has been paid to machine-generated version information. In other words, both previous approaches rely on the version information contained in or attached to the ontology. If such information is missing or incorrect, there is no way to automatically detect versioning of ontology concepts. In this study, we introduce a lightweight and fully automatic OV approach. Our approach advances by reusing existing OM systems and techniques for OV tasks, rather than creating a new OV framework from scratch. To the best of our knowledge, our work is the first to systematically analyse and utilise OM for OV.
@@ 3 OM4OV Pipeline
The OM task is to find an alignment $A$ with respect to a given similarity threshold $s\in[0,1]$, defined as [6]: $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},s\leq c<1\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$), subsumption ($\subseteq$), or other relations, and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$.
Similarly, an OV task can be formalised as finding two variants of an alignment $A_{match}$ and $A_{non-match}$ between ontology versions $O$ and $O^{\prime}$ with respect to a given similarity threshold $s\in[0,1]$. While OM may have different mapping relations, OV narrows down the task and focuses only on the equivalence relation.
Definition 3.2.
We classify four subset alignments produced in the OV process, namely $A_{remain}$ ($A\odot$), $A_{update}$ ($A\otimes$), $A_{add}$ ($A\oplus$), and $A_{delete}$ ($A\ominus$).
@@
The remain and update entities are actually matched entities between different ontology versions. The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions, assuming that the naming convention does not change between different ontology versions. If the naming convention changes, we will classify ontology entities that have changed names according to the systematic convention as update entities, by modelling their confidence $c$ as a high number more than $s$ but less than 1. The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions.
Definition 3.3.
$A\odot$ and $A\otimes$ are defined as:
$A\odot$ $=A_{match}(c=1)$ (2)
@@
$A\oplus$ and $A\ominus$ are defined as:
$A\oplus=\{e2\in O^{\prime}$ $\cap A_{non-match}\}=\{e2\in O^{\prime}\setminus A_{match}\}$ (3)
$A\ominus=\{e1\in O$ $\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\setminus A_{match}\}$
Given a gold standard reference (R) and a system-discovered alignment (A), OM typically measures performance using precision (Prec), recall (Rec), and the F1 score (F1). While precision measures matching correctness and recall measures matching completeness, the F1 score offers a harmonic mean to balance correctness and completeness.
Definition 3.5.
OV can reuse these measures, but they need to be extended into four sub-measures for add, delete, remain, and update performance. They are defined as:
$Prec\odot\otimes\oplus\ominus=\frac{|A\cap R|}{|A|}\quad Rec\odot\otimes\oplus\ominus=\frac{|A\cap R|}{|R|}\quad F_{1}\odot\otimes\oplus\ominus=\frac{2}{Prec^{-1}+Rec^{-1}}$ (4)
@@ 4.1 Dataset Construction
There is a dearth of lab-based datasets for the evaluation of OV. The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks; however, none of them includes benchmarks specifically designed for OV tasks. We propose an approach to constructing synthetic OV datasets from OM datasets. Figure 2 illustrates the generation of OM4OV datasets. Given a source ontology and target ontology with their reference alignment reference.xml, our approach is described in the following steps:
@@
The original OAEI datasets for OM provide two ontologies, $O_{s}$ and $O_{t}$. We choose one to be the intermediate ontology $O_{i}$. We retrieve all ontology entities from $O_{i}$.
There are four possible changes in OV tasks: remain, update, add, and delete. Each ontology entity in $O_{i}$ is randomly assigned one of these. For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$. Elements of update are necessarily mappings $(e,e^{\prime})$ (also written $e\rightarrow e^{\prime})$ where $e\neq e^{\prime}$ and $e^{\prime}$ is the updated entity name of $e$, but when we write $e\in\textit{update}$ we mean $(e,e^{\prime})\in\textit{update}$ for some $e^{\prime}$. By construction, we also have the four sets as pairwise disjoint.
For entities assigned to update, we need to generate the updated entity name. We should expect the new name to have a similar meaning to its original name. For example, the entity name “ConferenceVenuePlace” could be replaced with “Conference_hall” or “Conference_building”, but not the general names “Place” or “Location”. To achieve this goal, we retrieve all equivalent entities provided by reference.xml included in the original OAEI datasets and use them as a replacement synonym corpus. For entities whose names are unique identifiers or codes (and not textually meaningful names), we use their annotation properties (e.g. rdfs:label, rdfs:comment, skos:prefLabel, and skos:definition) instead. For entities that do not have synonyms in the generated list, we randomly re-assign them to remain, add, or delete.
Based on the entity assignments, we generate four corresponding versioning references, namely vr-remain.xml, vr-update.xml, vr-add.xml, and vr-delete.xml. For entities assigned to remain, no operation is required. For entities assigned to update, add, and delete, we will generate a corresponding update, add, and delete list.
@@
Let $e$ be an entity and $A$ be a mapping of the form $\{e_{1}\rightarrow e^{\prime},e_{2}\rightarrow e^{\prime\prime},\ldots e_{n}\rightarrow e^{{}^{\prime\prime}...^{\prime}}\}$. Then $map(e,M)$ is defined to be $e^{\prime}$ if there is an $e^{\prime}$ such that $e\rightarrow e^{\prime}\in A$ and to be $e$ otherwise. Now, $O^{\prime}$ = $O_{i}\setminus\{(s,p,o)|(s,p,o)\textit{ is a triple}\in O_{i}\textit{ and }s\in\textit{delete}\textit{ or }p\in\textit{delete}\textit{ or }o\in\textit{delete}\}\cup\{(s^{\prime},p^{\prime},o^{\prime})|(s,p,o)\textit{ is a triple}\in O_{i}$ and $s^{\prime}=map(s,\textit{update})\textit{ or }p^{\prime}=map(p,\textit{update})\textit{ or }o^{\prime}=map(o,\textit{update})\}$. That is, $O^{\prime}$ is constructed as $O_{i}$ without all the triples related to entities in the delete list and updated for all the triples related to entities in the update list.
We also incorporate the cross-reference mechanism into our OV testbed, $R_{or}$ and $R_{o^{\prime}r}$ are created according to the following rules: (1) $R_{or}$ is the original reference.xml that removes all the mappings related to add entities. (2) $R_{o^{\prime}r}$ is the original reference.xml that removes all the mappings related to delete entities and updates all the mappings related to update entities. Unlike the original OAEI datasets used for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed. This suits the dynamic nature of OV, where the changes vary between different versions. For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks. Table 1 lists the detailed information of the selected OAEI track used for the OV testbed. The current version of the OV testbed contains a total of 12 distinct ontologies from three different OAEI tracks. The anatomy track contains two large ontologies, while the MSE track has three medium ontologies, and the conference track has 7 small ontologies. Figure 3 shows the number of entities in each track.

### v4

@@ Abstract
Due to the dynamic nature of the Semantic Web, version control is necessary to capture time-varying information, particularly for widely used ontologies. Despite the long-standing recognition of ontology versioning (OV) as a crucial component for efficient ontology management, the growing size of ontologies and accumulating errors caused by manual labour overwhelm current OV approaches. In this paper, we propose a fresh approach to performing OV using existing ontology matching (OM) techniques and systems. We introduce a unified OM4OV pipeline. From an OM perspective, we reconstruct a new task formulation and measurements for OV tasks. Building upon the prior alignment(s) from OM, we propose a pipeline optimisation method called the cross-reference (CR) mechanism to enhance overall OV performance. We experimentally validate the OM4OV pipeline and the cross-reference mechanism in an OV testbed originating from the Ontology Alignment Evaluation Initiative (OAEI) datasets. We also discuss insights into OM used for OV tasks, where some apparent false mappings detected by OV systems are not actually untrue.
@@ 1 Introduction
Ontologies serve as the backbone of the semantic web, providing formal descriptions of shared concepts across various applications [8]. An ontology is not static, and the need for version control arises with its origination. While web data is dynamic, any ontology used must undergo periodic revisions to keep pace with the growth in domain knowledge, modifications to application adaptation, or corrections to the shared conceptualisation [13]. For example, it is unrealistic to expect ontologies created in the 1990s to contain concepts such as “touchscreen”, “fingerprint sensor”, or “WiFi antenna” [9]. This may cause undesirable deficiencies in artefacts that conform to or reuse the ontology that is being changed, leading to severe non-compliance and incompatibility issues in downstream tasks.
@@
Ontology versioning (OV) aims to distinguish and recognise changes between different ontology versions. By doing so, data that conforms to the changed ontology, other ontologies that reuse the changed ontology, or software that uses the changed ontology can apply the correct changes correspondingly [13]. While various methods for OV have been developed, one approach is to extend the ontology itself with internal version information. An ontology can be issued with a unique identifier (e.g. IRI) or a specific version number (e.g. owl:versionInfo) to be distinguished from other versions; each ontology entity may have a new annotation to record its current status (e.g. owl:DeprecatedClass and owl:DeprecatedProperty); or every triple could be extended with a 4th dimension populated with a triple timestamp, similar to the RDF-star schema. Alternatively, version information can be recorded in change logs. Change logs can take the form of free-text notebooks, an extensible markup language (XML) document, or a knowledge graph (KG). However, maintaining version information in the ontology can be time-consuming and labour-intensive. Either extending the current schema or using change logs requires consistent updating over time. In most cases, this process is hand-crafted by the ontology engineer or requires human intervention (e.g. pre-defining a schema or creating a template for the change log). A manual process is more likely to make mistakes and fail to propagate changes to dependent artefacts. Also, in the real world, there is no guarantee that the ontology itself contains the complete version information or that it has a separate change log. In such cases, current approaches have limited capabilities in detecting incorrect version information or identifying missing version information.
@@
In this study, we investigate a lightweight and fully automatic version control approach for ontologies. We observe that the nature of OV is very similar to that of ontology matching (OM). Both are introduced to address interoperability between ontologies, with ontology entities serving as inputs. While OM is a well-studied problem [17], a unified pipeline that can extend and reuse existing OM techniques and systems for OV tasks would be highly desirable. However, OM techniques and systems cannot be directly applied to OV tasks because OM and OV have some significant differences. The OM input is two distinct ontologies, while the OV input is expected to be two different versions of a single ontology. The output of OM is a set of entity mappings, whereas the output of OV comprises two sets: changed entities and unchanged entities. OM concentrates on similarities between two entities, while OV addresses differences between versions of one entity. OM determines matched entities between two different ontologies, while OV distinguishes add, delete, remain, and update entities between different versions of one ontology. To address these challenges, our goal is to investigate the complementary relationship between these two tasks and to develop a unified approach to shifting OV tasks towards OM tasks, thereby leveraging the deep research legacy of OM to assist with OV. Specifically, our key contributions include:
@@
  • We argue that some false mappings detected by OV systems are not actually untrue, but can be caused by a flawed choice of ontology design, an ambiguous “equivalent” relationship, or an inappropriate setting of a similarity threshold.
The remainder of the paper is organised as follows. Section 2 reviews the literature on OV. Section 3 introduces our novel OM4OV pipeline, while Section 4 evaluates the proposed pipeline. We propose the cross-reference mechanism as a pipeline optimisation method and experimentally evaluate its performance in Section 5. Section 6 and 7 discuss our implications and current limitations. Section 8 concludes the paper.
@@ 2 Related Work
One option is to include version information within the ontology. The Simple HTML Ontology Extensions (SHOE) [11] uses the tag BACKWARD-COMPATIBLE-WITH to record version information. The authors in [12] argue that a carefully-managed version numbering system embedded in the URI of the ontology (and therefore the fully expanded name of entities defined in the ontology) can minimise the impact of adopting updated versions because unchanged entities will be unaffected in practice. These approaches have largely been adopted by the later OWL Web Ontology Language [4], where a set of annotation properties related to version information is defined. These include owl:versionInfo and owl:priorVersion to describe the version number of the ontology, owl:backwardCompatibleWith and owl:incompatibleWith to specify an entity’s compatible or incompatible corresponding entity in the previous version, and owl:DeprecatedClass and owl:DeprecatedProperty to declare deprecated entities. Later, the ontology language $\tau$OWL [22] was introduced to extend the OWL triple schema to quadruples to represent the versioning of concepts within an ontology. This idea is now incorporated in the new proposals for RDF 1.2, which allow time-varying information to be deduced from a temporal dimension within the quadruple [20].
@@
While current approaches simply record human-generated version information, less attention has been paid to machine-generated version information. In other words, both previous approaches rely on the version information contained in or attached to the ontology. If such information is missing or incorrect, there is no way to automatically detect versioning of ontology concepts. In this study, we introduce a lightweight and fully automatic OV approach. Our approach advances by reusing existing OM systems and techniques for OV tasks, rather than creating a new OV framework from scratch. To the best of our knowledge, our work is the first to systematically analyse and utilise OM for OV. The output of our novel automatic OV detection could be easily recorded by any of the current approaches presented above.
@@ 3 OM4OV Pipeline
The OM task is to find an alignment $A$ with respect to a given similarity threshold $s\in[0,1]$, defined as [6]: $A=\{(e1,e2,r,c)|e1\in O_{s},e2\in O_{t},s\leq c<1\}$, where $e1$ and $e2$ are ontology entities in $O_{s}$ and $O_{t}$ respectively, $r$ states the relation between $e1$ and $e2$ which can be equivalence ($\equiv$), subsumption ($\subseteq$), or other relations, and $c\in[0,1]$ is the level of confidence for the match $(e1,r,e2)$. Similarly, an OV task can be formalised as finding two variants of an alignment $A_{match}$ and $A_{non-match}$ between ontology versions $O$ and $O^{\prime}$ with respect to a given similarity threshold $s\in[0,1]$. While OM may have different mapping relations, OV narrows down the task and focuses only on the equivalence relation.
@@
An OV task is defined as:
@@
We classify four subset alignments produced in the OV process, namely $A_{remain}$ ($A\odot$), $A_{update}$ ($A\otimes$), $A_{add}$ ($A\oplus$), and $A_{delete}$ ($A\ominus$). The remain and update entities are actually matched entities between different ontology versions. The remain entities can be considered to be the exact (trivial) equivalent string matches between different ontology versions, assuming that the naming convention does not change between different ontology versions. If the naming convention changes, we will classify ontology entities that have changed names according to the systematic convention as update entities, by modelling their confidence $c$ as a high number more than $s$ but less than 1. The update entities are, in general, the closest (non-trivial) near-equivalent matches between different ontology versions.
Definition 3.1.1.
$A_{match}$, $A\odot$ and $A\otimes$ are defined as:
$A_{match}=A\odot\cup A\otimes$ (2)
$A\odot=A_{match}(c=1)$
@@
Definition 3.1.2.
$A_{non-match}$, $A\oplus$, and $A\ominus$ are defined as:
$A_{non-match}=A\oplus\cup A\ominus$ (3)
$A\oplus=\{e2\in O^{\prime}\cap A_{non-match}\}=\{e2\in O^{\prime}\setminus A_{match}\}$
$A\ominus=\{e1\in O\phantom{{}^{\prime}}\cap A_{non-match}\}=\{e1\in O\phantom{{}^{\prime}}\setminus A_{match}\}$
Given a gold standard reference (R) and a system-discovered alignment (A), OM typically measures performance using precision (Prec), recall (Rec), and the F1 score (F1). While precision measures matching correctness and recall measures matching completeness, the F1 score offers a harmonic mean to balance correctness and completeness. OV can reuse these measures, but they need to be extended into four sub-measures for add, delete, remain, and update performance.
@@
The measures of OV task are defined as:
$Prec\odot|\otimes|\oplus|\ominus=\frac{|A\cap R|}{|A|}\quad Rec\odot|\otimes|\oplus|\ominus=\frac{|A\cap R|}{|R|}\quad F_{1}\odot|\otimes|\oplus|\ominus=\frac{2}{Prec^{-1}+Rec^{-1}}$ (4)
@@ 4.1 Dataset Construction
There is a dearth of benchmark datasets for the evaluation of OV. The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks; but none are specifically designed for OV tasks. We propose an approach to constructing synthetic OV datasets from OM datasets. Figure 2 illustrates the generation of OM4OV datasets. Given a source ontology and a target ontology with their reference alignment reference.xml, our approach is described in the following steps:
@@
The original OAEI datasets for OM provide two ontologies, the source $O_{s}$ and the target $O_{t}$. We choose one to be the intermediate ontology $O_{i}$. We retrieve all ontology entities from $O_{i}$.
There are four possible entity changes in OV tasks: remain, update, add, and delete. Each ontology entity in $O_{i}$ is randomly assigned to one of these. For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$. Elements of update are necessarily mappings $(e,e^{\prime})$ (also written $e\rightarrow e^{\prime})$ where $e\neq e^{\prime}$ and $e^{\prime}$ is the updated entity name of $e$, but when we write $e\in\textit{update}$ we mean $(e,e^{\prime})\in\textit{update}$ for some $e^{\prime}$. By construction, we also have the four sets as pairwise disjoint.
For entities assigned to update, we need to generate a new name for the updated entity. We should expect the new name to have a similar meaning to its original name. For example, the entity name “ConferenceVenuePlace” could be replaced with “Conference_hall” or “Conference_building”, but not the general names “Place” or “Location”. To achieve this goal, we retrieve all equivalent entities provided by reference.xml included in the original OAEI datasets and use them as a replacement synonym corpus. For entities whose names are unique identifiers or codes (and not textually meaningful names), we use their annotation properties (e.g. rdfs:label, rdfs:comment, skos:prefLabel, and skos:definition) instead of the code. For entities in update that do not have synonyms in the corpus, we randomly re-assign them to remain, add, or delete.
Based on the entity assignments, we generate four corresponding versioning references, namely vr-remain.xml, vr-update.xml, vr-add.xml, and vr-delete.xml. For entities assigned to update, add, and delete, we will generate a corresponding update, add, and delete list. For entities assigned to remain, no operation is required.
@@
Let $e$ be an entity and $M$ be a mapping of the form $\{e_{1}\rightarrow e^{\prime},e_{2}\rightarrow e^{\prime\prime},\ldots e_{n}\rightarrow e^{{}^{\prime\prime}...^{\prime}}\}$. Then $map(e,M)$ is defined to be $e^{\prime}$ if there is an $e^{\prime}$ such that $e\rightarrow e^{\prime}\in M$ and to be $e$ otherwise. Now, $O^{\prime}$ = $O_{i}\setminus\{(s,p,o)|(s,p,o)\textit{ is a triple}\in O_{i}\textit{ and }s\in\textit{delete}\textit{ or }p\in\textit{delete}\textit{ or }o\in\textit{delete}\}\cup\{(s^{\prime},p^{\prime},o^{\prime})|(s,p,o)\textit{ is a triple}\in O_{i}$ and $s^{\prime}=map(s,\textit{update})\textit{ or }p^{\prime}=map(p,\textit{update})\textit{ or }o^{\prime}=map(o,\textit{update})\}$. That is, $O^{\prime}$ is constructed as $O_{i}$ without all the triples related to entities in the delete list and updated for all the triples related to entities in the update list.
We also incorporate the cross-reference mechanism into our OV testbed, $R_{or}$ and $R_{o^{\prime}r}$ are created according to the following rules: (1) $R_{or}$ is the original reference.xml that removes all the mappings related to add entities. (2) $R_{o^{\prime}r}$ is the original reference.xml that removes all the mappings related to delete entities and updates all the mappings related to update entities. Unlike the original OAEI datasets used for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed. This suits the dynamic nature of OV, where the changes vary between different versions. For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks. For reproducibility, we use a fixed seed to control the randomness in steps (3) and (4).
4.2 Experiment Setup
Table 1 lists the detailed information of the selected OAEI track used for the OV testbed. The current version of the OV testbed contains a total of 12 distinct ontologies from three different OAEI tracks. The anatomy track contains two large ontologies, while the MSE track has three medium ontologies, and the conference track has 7 small ontologies. Figure 3 shows the number of entities in each track. The fixed seed for the OV testbed is set to 42.